MILLIMETER-WAVE SPECTROSCOPY OF NCS:AN ANALYSIS OF VIBRONIC INTERACTION

\underline{TAKAYOSHI AMANO}, Institute for Astrophysics and Planetary Sciences, Ibaraki University, Mito 310, Japan..

The ground electronic state of NCS radical is a ^{2}Π_{i} state and is isovalent with NCO. While NCO is one of the free radicals that have been studied very extensively in the microwave, millimeter-wave and visible-UV regions, high-resolution investigation of NCS has been relatively sparse. We studied the millimeter-wave spectrum of this radical in the ground vibrational state several years ago, and determined precise set of the molecular constants^{a}. Since then we extended our measurements to the excited vibrational states, v_{2} = 1 state in particular. The spectrum was found to be complicated due to large Renner-Teller interaction. The Δ and Σ vibronic states are fitted separately using effective Hamiltonians. A more complete global fit including the v_{2} = 2 vibronic states is in progress and will be discussed.